Determine the minimum number of bits you would need for a memory address if memory could have up to 68GBytes. Assume byte addressing and explain how do you arrived at answer??

A).Since there is no calculator available, solve it by estimation method.

B). If a calculator is available, show me alternative approach to solve the problem.
Answer #1

68 GBs is same as 68 * 2^30 bytes

for byte addressing, each byte has an adsress

64*2^30 < 68*2^30 < 128*2^30

64*2^30 = 2^36

128*2^30 = 2^37

this means 68GB of memory bytes need 37 bits to store address.

With calculator

68 GB is 68*1024*1024*1024 bytes

to find nearest power of 2, take log2(68*1024*1024*1024) = 36.08 > 36 so we need 37 bits